# Early stopping for conjugate gradients in statistical inverse problems

### v1

@@ Abstract
The error analysis of CG under statistical noise is subtle due to its nonlinear dependence on the observations. We provide an explicit error decomposition into two terms, which shares important properties of the classical bias-variance decomposition. Together with a continuous interpolation between CG iterates, this paves the way for a comprehensive error analysis of early stopping. In particular, a general oracle-type inequality is proved for the prediction error at $\tau$. For bounding the reconstruction error, a more refined probabilistic analysis, based on concentration of self-normalised Gaussian processes, is developed. The methodology also provides some new insights into early stopping for CG in deterministic inverse problems. A numerical study for standard examples shows good results in practice for early stopping at $\tau$.
@@ 1 Introduction
In the following Section 2, the setting is introduced. Section 3 explains the main ideas and results of the paper. Afterwards, we progress in detail. Results for the interpolated CG iterates are presented in Section 4, and the prediction error is analysed in Section 5. Section 6 derives our stopping criterion and proves an oracle-type inequality for its prediction error. The results are transferred in Section 7 to the reconstruction error. The numerical results for some standard examples in Section 8 are in good agreement with the theory. Technical results and proofs are delegated to Section 9.
@@ 2 The setting
$\lVert f\rVert_{\mu}\coloneqq\Big{(}\sum_{i=1}^{D}\lambda_{i}^{-4\mu}f_{i}^{2}\Big{)}^{1/2}\leqslant R$
for $\mu>0$, $R\geqslant 1$. Note that the assumption $\lVert f\rVert_{\mu}\leqslant R$ on $f$ corresponds to $\lVert g\rVert_{\mu+1/2}\leqslant R$ on $g=Af$.
@@
where $p\geqslant 0$ is fixed and $0<c_{A}\leqslant C_{A}<\infty$ are numerical constants.
Throughout the paper, (in-)equalities up to absolute constants are denoted by $\cong$’, $\lesssim$’, $\gtrsim$’. Further dependencies on the constants will be highlighted by additional indices. Bounds of the type $B_{1}\lesssim_{c_{A}}B_{2}$, $B_{3}\lesssim_{C_{A}}B_{4}$ indicate that they hold true already under a one-sided condition, i.e. for $C_{A}=\infty$ and $c_{A}=0$, respectively.
@@
Let us introduce some further notation. Define
@@ 3 Overview of main results
$r_{k}\coloneqq\operatorname*{argmin}_{p_{k}\in\operatorname{Pol}_{k,1}}\lVert p_{k}(AA^{\top})Y\rVert^{2}.$
In the case $k=d$, $r_{d}$ maps all squared nonzero singular values of $A$ to zero implying $\widehat{f}_{d}=A^{\dagger}Y$. Thus, the CG estimator at iteration $d$ is a minimum-norm solution of (2.1) and $d$ is the maximal iteration index.
One new idea is to allow also for noninteger iteration indices $t\in[0,d]$ by interpolating linearly between the residual polynomials and thus also between the estimators. The time-continuous residual polynomials have useful properties, which are discussed in Section 4 in detail.
Since the estimator $\widehat{g}_{t}$ depends nonlinearly on the observation $Y$, we cannot calculate a standard bias-variance decomposition, but need to analyse the prediction error for any realisation of $\xi$. In Proposition 5.2, we are able to decompose the error nonasymptotically into terms that share important properties of stochastic and approximation errors, see Lemma 5.3 for a detailed analysis.
Corollary 3.2.
For $t\in[0,d]$, the prediction error satisfies
$\lVert A(\widehat{f}_{t}-f)\rVert^{2}\leqslant 2(A_{t,\lambda}+S_{t,\lambda}),$
where we have for the stochastic error term $S_{t,\lambda}$ that
  1. (a)
    $S_{t,\lambda}=0$ for $\xi=0$,
  2. (b)
    $S_{0,\lambda}=0$ and $S_{d,\lambda}=\lVert\xi\rVert^{2}$,
  3. (c)
    $S_{t,\lambda}$ is nondecreasing in $t$,
and for the approximation error term $A_{t,\lambda}$ that
  1. (a)
    $A_{t,\lambda}\leqslant 0$ if $Af=0$,
  2. (b)
    $A_{0,\lambda}=\lVert Af\rVert^{2}$ and $A_{d,\lambda}=0$,
  3. (c)
    $A_{t,\lambda}$ has a nonnegative upper bound given in 5.3(h), which is nonincreasing in $t$.
Both error terms, defined in Definition 5.1, depend on the observation through the residual polynomial. In addition, the stochastic error depends on the noise $\xi$ and the approximation error on the signal $g=Af$. Therefore, they are unknown to the user, but they play a key role in the nonasymptotic analysis of early stopping. Given the monotonicity properties of $S_{t,\lambda}$ and the upper bound of $A_{t,\lambda}$, it is natural to consider the following sequential oracle index.
@@
Due to the continuity of the interpolations $A_{t,\lambda}$ and $S_{t,\lambda}$ in $t$ and $A_{0,\lambda}\geqslant S_{0,\lambda}$, $A_{d,\lambda}\leqslant S_{d,\lambda}$, the approximation and stochastic error are indeed balanced at $\tau_{\mathfrak{b}}$, meaning that $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$.
@@
$A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda}=2S_{\tau_{\mathfrak{b}},\lambda}\leqslant 2\inf_{t\in[0,d]}(A_{t,\lambda}\vee S_{t,\lambda})\leqslant\inf_{t\in[0,d]}2(A_{t,\lambda}+S_{t,\lambda}).$
@@
$\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)\coloneqq R^{2/(4\mu p+2p+1)}\delta^{(8\mu p+4p)/(4\mu p+2p+1)}$
@@
From an asymptotic point of view, where the dimension $D=D(\delta)\to\infty$ as $\delta\to 0$, this rate is minimax optimal if $D(\delta)$ grows faster than $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$ as shown in Proposition 5.11. The proof of the upper bound is very general and allows extensions to other source conditions and noise specifications (see Remark 5.10). Also it enables us to show close parallels to Showalter’s method, i.e. standard gradient flow (see Remark 5.13).
For deterministic noise of level $\delta$, the prediction error at the least-squares solution (i.e. at the last iteration $\widehat{f}_{d}$) is $\delta^{2}$ and no regularisation is required. Under (GN), however, this becomes $\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{d}-f)\rVert^{2}]=D\delta^{2}$ and regularisation by early stopping is essential for larger dimensions $D$ exactly as in nonparametric regression. It is the first time that the CGNE iteration path is shown to achieve the optimal prediction error rate under (GN) at some iterate.
@@
Under (DN), setting $\kappa\coloneqq c^{2}\delta^{2}$ for an a-priori fixed constant $c>1$, this coincides with the well-known discrepancy principle [8, Sections 4.3 and 7.3]. The balanced oracle can also be expressed in terms of the residual norm, and in Remark 6.2, the relation between the stopping indices $\tau$ and $\tau_{\mathfrak{b}}$ is discussed.
@@
Note that we do not have a classical oracle inequality in terms of the oracle error $\inf_{t\in[0,d]}\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{t}-f)\rVert^{2}]$, since the decomposition into the two error terms in Corollary 3.2 gives only an upper bound. The dimension-dependent error term $\delta^{2}\sqrt{D}$ is due to the variability of the Gaussian noise $\xi$ and unavoidable in the linear case of truncated SVD estimation [3]. Therefore, the early stopped estimator $\widehat{g}_{\tau}$ has the same error rate as $\widehat{g}_{\tau_{\mathfrak{b}}}$ if the remainder term is small enough. In this case of a-posteriori error control, we say that we can adapt to the balanced oracle. In practice, the noise level $\delta$ needs to be estimated, and the threshold $\kappa$ is allowed to deviate from $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]=\delta^{2}D$.
@@
and the optimal rate of convergence is attained. The early stopping rule $\tau$ is minimax adaptive for the prediction error over Sobolev-type ellipsoids for all regularities $\mu>0$ with $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$ (see Corollary 6.10).
In Section 7, we transfer the results for the prediction error to the reconstruction error, i.e. the risk of the CG estimator. Similarly to the prediction error, the analysis of the reconstruction error relies on an upper bound given in Proposition 7.1, which depends on the error terms and the residual polynomial.
@@
$\sup_{f:\,\lVert f\rVert_{\mu}\leqslant R}\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{f}_{\tau_{\mathfrak{b}}}-f^{\dagger}\rVert^{2}\big{]}\lesssim_{\mu,p,C_{A}}\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)\coloneqq R^{(4p+2)/(4\mu p+2p+1)}\delta^{8\mu p/(4\mu p+2p+1)}.$
@@ 4 The interpolated conjugate gradient method
The interpolation follows also naturally by defining $r_{t}$ via penalised minimisation. Let
$r_{k+1,\lambda}\coloneqq\operatorname*{argmin}_{p_{k+1}\in\operatorname{Pol}_{k+1,1}}\Big{(}\lVert\mathsf{p_{k+1}}\rVert_{Y}^{2}+\lambda\alpha_{k+1}(p_{k+1})^{2}\Big{)},$
where $\alpha_{k+1}(p)\in{\mathbb{R}}$ denotes the coefficient in front of $x^{k+1}$ for a polynomial $p(x)$ and $\lambda>0$ is a penalty parameter. Note that $p_{k+1}-r_{k}-\alpha(r_{k+1}-r_{k})$ lies in $\operatorname{Pol}_{k,0}$ for $\alpha=\alpha_{k+1}(p_{k+1})/\alpha_{k+1}(r_{k+1})$ and $r_{k},r_{k+1}\perp_{Y}\operatorname{Pol}_{k,0}$ by Lemma 9.3. By rewriting $p_{k+1}=r_{k}+\alpha(r_{k+1}-r_{k})+q_{k}$ with $q_{k}\in\operatorname{Pol}_{k,0}$ and applying the Pythagorean theorem, the minimisation yields a polynomial $r_{k+1,\lambda}=r_{k}+\alpha_{\lambda}(r_{k+1}-r_{k})$ with
$\alpha_{\lambda}\coloneqq{}$ $\operatorname*{argmin}_{\alpha\in{\mathbb{R}}}\Big{(}\lVert\mathsf{r_{k}}+\alpha(\mathsf{r_{k+1}-r_{k}})\rVert_{Y}^{2}+\lambda\alpha_{k+1}(r_{k+1})^{2}\alpha^{2}\Big{)}$
[…]
where we used that $\langle{\mathsf{r_{k+1}-r_{k}}},{\mathsf{r_{k+1}}}\rangle_{Y}=0$ by Lemma 9.3. The numerator is just $R_{k}^{2}-R_{k+1}^{2}$, which is strictly positive for $k+1\leqslant d$ so that $\alpha_{\lambda}\in(0,1]$ holds with $\alpha_{\lambda}\downarrow 0$ for $\lambda\uparrow\infty$ and $\alpha_{\lambda}=1$ for $\lambda=0$. This way we recover $r_{t}$ for $t=k+\alpha$, $k=0,\dots,d-1$, $\alpha\in(0,1]$, as $r_{k+1,\lambda}$ for a suitable penalisation parameter $\lambda>0$.
We have extended Definition 3.1 for $t=1,\dots,d$, keeping the starting values $\widehat{g}_{0}=\widehat{f}_{0}=0$ and the relation $A\widehat{f}_{t}=\widehat{g}_{t}$ for $t\in[0,d]$. Since there are $d$ distinct nonzero singular values of $A$, the $k$-th residual polynomial has degree $k$ for $k=0,\dots,d$. Therefore, $r_{t}$ for $t\in(k,k+1]$ has degree $k+1$ and is an element of $\operatorname{Pol}_{k+1,1}$. The squared residual norms interpolate nonlinearly:
@@
Lemma 4.3.
For $t,s\in[0,d]$, the squared residual norm satisfies
@@
and set $r_{0,<}\coloneqq r_{0}\equiv 1$, $r_{0,>}\equiv 0$. If $x_{1,t}$ is clear from the context, we also write $f_{<}(x)\coloneqq f(x){\bf 1}(x<x_{1,t})$, $\lVert\mathsf{f}\rVert_{Y,<}\coloneqq\lVert\mathsf{f}_{<}\rVert_{Y}$ and $\langle{\mathsf{f}},{\mathsf{g}}\rangle_{Y,<}\coloneqq\langle{\mathsf{f_{<}}},{\mathsf{g_{<}}}\rangle_{Y}$ for any functions $f$ and $g$, analogously for ‘$>$’.
For integer iteration indices, the residual polynomials possess several well-known properties, which are collected in Lemma 9.5. These carry over to the interpolated analogues. We state in the following lemma the most important properties.
@@
    For $k=1,\dots,d-1$, $r_{k}$ has $k$ real simple zeros $0<x_{1,k}<\dots<x_{k,k}<\lVert AA^{\top}\rVert_{\mathrm{spec}}$, where $\lVert\cdot\rVert_{\mathrm{spec}}$ denotes the spectral norm on ${\mathbb{R}}^{D\times D}$. For $k=d$, the zeros are the $d$ distinct squared nonzero singular values $0<\tilde{\lambda}_{d}^{2}<\dots<\tilde{\lambda}_{1}^{2}=\lVert AA^{\top}\rVert_{\mathrm{spec}}$.
@@
    For fixed $x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}]$, $t\mapsto r_{t,<}(x)$ is nonincreasing on $[0,d]$.
@@
Using the upper bound, we get directly the following control for the approximation error $A_{t,\lambda}$ defined below.

### v2

@@ Abstract
The error analysis of CG under statistical noise is subtle due to its nonlinear dependence on the observations. We provide an explicit error decomposition and identify two terms in the prediction error, which share important properties of classical bias and variance terms. Together with a continuous interpolation between CG iterates, this paves the way for a comprehensive error analysis of early stopping. In particular, a general oracle-type inequality is proved for the prediction error at $\tau$. For bounding the reconstruction error, a more refined probabilistic analysis, based on concentration of self-normalised Gaussian processes, is developed. The methodology also provides some new insights into early stopping for CG in deterministic inverse problems. A numerical study for standard examples shows good results in practice for early stopping at $\tau$.
@@ 1 Introduction
In the following Section 2, the setting is introduced. Section 3 explains the main ideas and results of the paper. Afterwards, we progress in detail. Results for the interpolated CG iterates are presented in Section 4, and the prediction error is analysed in Section 5. Section 6 derives our stopping criterion and proves an oracle-type inequality for its prediction error. The results are transferred to the reconstruction error in Section 7. The practical performance for some standard examples in Section 8 is in good agreement with the theory. Technical results and proofs are delegated to Section 9.
@@ 2 The setting
$\lVert f\rVert_{\mu}\coloneqq\Big{(}\sum_{i=1}^{D}\lambda_{i}^{-4\mu}f_{i}^{2}\Big{)}^{1/2}\leqslant R$ (2.2)
for $\mu>0$, $R\geqslant 1$. Note that $f^{\dagger}_{i}\coloneqq\langle{f^{\dagger}},{v_{i}}\rangle=f_{i}$ for $i=1,\dots,D$ such that $\lVert f^{\dagger}\rVert_{\mu}=\lVert f\rVert_{\mu}$ and the respective source conditions agree. The assumption $\lVert f\rVert_{\mu}\leqslant R$ on $f$ corresponds to $\lVert g\rVert_{\mu+1/2}\leqslant R$ on $g=Af$.
@@
where $p\geqslant 0$ is fixed and $0\leqslant c_{A}\leqslant C_{A}\leqslant\infty$. If $C_{A}=\infty$ or $c_{A}=0$, we explicitly write (PSD($p$,$c_{A}$$\infty$)) or (PSD($p$,00,$C_{A}$)), respectively. Throughout the paper, (in-)equalities up to absolute constants are denoted by $\cong$’, $\lesssim$’, $\gtrsim$’. Further dependencies on the constants will be highlighted by additional indices. Bounds of the type $B_{1}\lesssim_{c_{A}}B_{2}$, $B_{3}\lesssim_{C_{A}}B_{4}$ indicate that they hold true already under a one-sided condition, i.e. for $C_{A}=\infty$ and $c_{A}=0$, respectively.
@@
Let us comment on the relationship between our a-priori discrete theory and the potentially underlying continuous inverse problem. The CG algorithm is always applied to a discretised problem, and our analysis will provide nonasymptotic results, which are directly interpretable. In order to assess the size of different terms, we then also adopt a classical asymptotic view, where the noise level $\delta$ tends to zero and simultaneously the dimensions $D=D(\delta),P=P(\delta)$ tend to infinity. As $D(\delta),P(\delta)\to\infty$, we may approach a continuous inverse problem in terms of an operator $A$ and a function $f$ such that the source condition (2.2) and the spectral decay assumption PSD($p$,$c_{A}$,$C_{A}$) may a-priori vary with the discretisation $D$, and we refer to Remark 5.12 below for concrete consequences.
@@
Let us introduce some further notation. Applying standard functional calculus for symmetric matrices, we define
@@ 3 Overview of main results
$r_{k}\coloneqq\operatorname*{argmin}_{p_{k}\in\operatorname{Pol}_{k,1}}\lVert p_{k}(AA^{\top})Y\rVert^{2}=\operatorname*{argmin}_{p_{k}\in\operatorname{Pol}_{k,1}}\lVert\mathsf{p_{k}}\rVert_{Y}^{2}.$
In the case $k=d$, by 4.5(a), $r_{d}$ maps all squared nonzero singular values of $A$ to zero implying $\widehat{f}_{d}=A^{\dagger}Y$. Thus, the CG estimator at iteration $d$ is a minimum-norm solution of (2.1) and $d$ is the maximal iteration index.
One new idea is to allow also for noninteger iteration indices $t\in[0,d]$ by interpolating linearly between the residual polynomials and thus also between the estimators (see Definition 4.1). The time-continuous residual polynomials have useful properties, which are discussed in Section 4 in detail.
We start by considering the prediction error $\lVert A(\widehat{f}_{t}-f)\rVert^{2}=\lVert\widehat{g}_{t}-g\rVert^{2}$, i.e. the error of the prediction estimator $\widehat{g}_{t}=A\widehat{f}_{t}$. The prediction error is the intrinsic metric for the minimisation property of CG as well as for a-posteriori error control based on the residuals. Classically, interpolation often allows to transfer prediction results to the reconstruction (or estimation) error $\lVert\widehat{f}_{t}-f^{\dagger}\rVert^{2}$ [8, Section 4.3]. While in deterministic situations the prediction error is sometimes also analysed for its own sake [12, Section 4.5], optimal regularisation is essential for prediction in statistical inverse problems and there is a strong genuine interest in the analysis of the prediction error.
Since the estimator $\widehat{g}_{t}$ depends nonlinearly on the observation $Y$, we cannot calculate a standard bias-variance decomposition, but need to analyse the prediction error for any realisation of $\xi$. In Proposition 5.2, we are able to decompose the error nonasymptotically and obtain for $t\in[0,d]$ the bound
$\lVert A(\widehat{f}_{t}-f)\rVert^{2}\leqslant 2(A_{t,\lambda}+S_{t,\lambda}),$ (3.1)
where $S_{t,\lambda}$ and $A_{t,\lambda}$ are defined in Definition 5.1. The subscript $\lambda$ indicates that the terms correspond to the prediction error, which involves the singular values of $A$, in line with the notation in [3, 4]. As shown in Lemma 5.3, $S_{t,\lambda}$ and $A_{t,\lambda}$ share important properties of stochastic and approximation errors. In particular, the stochastic error term $S_{t,\lambda}$ is continuously increasing from $S_{0,\lambda}=0$ to $S_{d,\lambda}=\lVert\xi\rVert^{2}$. The approximation error term $A_{t,\lambda}$ is also continuous and has a relatively tight upper bound, which is decreasing from $A_{0,\lambda}=\lVert Af\rVert^{2}$ to $A_{d,\lambda}=0$. These properties are the key for our analysis.
Given the monotonicity properties of $S_{t,\lambda}$ and of the upper bound of $A_{t,\lambda}$, it is natural to consider the following sequential oracle index.
@@
Due to the continuity of the interpolations $A_{t,\lambda}$ and $S_{t,\lambda}$ in $t$ and $A_{0,\lambda}\geqslant S_{0,\lambda}$, $A_{d,\lambda}\leqslant S_{d,\lambda}$, the index $\tau_{\mathfrak{b}}\in[0,d]$ exists and the errors are indeed balanced at $\tau_{\mathfrak{b}}$, meaning that $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$.
@@
$A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda}=2S_{\tau_{\mathfrak{b}},\lambda}\leqslant 2\inf_{t\in[0,d]}(A_{t,\lambda}\vee S_{t,\lambda})\leqslant 2\inf_{t\in[0,d]}(A_{t,\lambda}+S_{t,\lambda}).$
@@
$\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)\coloneqq R^{2/(4\mu p+2p+1)}\delta^{(8\mu p+4p)/(4\mu p+2p+1)}$ (3.2)
@@
From an asymptotic point of view, where the dimension $D=D(\delta)\to\infty$ as $\delta\to 0$, this rate is minimax optimal if $D(\delta)$ grows faster than $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$ as shown in Proposition 5.11. The proof of the upper bound is very general and allows extensions to other source conditions and noise specifications, see Remark 5.10. It exhibits close parallels to Showalter’s method or standard gradient flow, see Remark 5.13. This is the first time that the CGNE iteration path is shown to achieve the optimal prediction error rate under (GN) at some iterate.
For deterministic noise of level $\delta$, the prediction error at the least-squares solution (i.e. at the last iteration $\widehat{f}_{d}$) is $\delta^{2}$. Under (GN), however, this becomes $\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{d}-f)\rVert^{2}]=\delta^{2}D$ and regularisation by early stopping is essential for larger dimensions $D$ exactly as in nonparametric regression. Still, in case of small $D\ll(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$, the estimation problem becomes in a statistical sense parametric and no regularisation is needed. The rate $\delta^{2}D$, attained by $\widehat{f}_{d}$, is then minimax optimal for the prediction error, see Remark 5.12.
@@
Under (DN), setting $\kappa\coloneqq c^{2}\delta^{2}$ for an a-priori fixed constant $c>1$, this coincides with the well-known discrepancy principle [8, Sections 4.3 and 7.3]. Note that by the linear interpolation, $R_{t}^{2}$ is continuous in $t$ with $R_{0}^{2}=\lVert Y\rVert^{2}$ and $R_{d}^{2}=0$. Hence, $\tau\in[0,d]$ exists and we have the identity $R_{\tau}^{2}=\kappa$, provided $\kappa\leqslant\lVert Y\rVert^{2}$. The balanced oracle can also be expressed in terms of the residual norm, and in Remark 6.2, the relation between the stopping indices $\tau$ and $\tau_{\mathfrak{b}}$ is clarified.
@@
Note that we do not have a classical oracle inequality in terms of the oracle error $\inf_{t\in[0,d]}\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{t}-f)\rVert^{2}]$, since the decomposition into the two error terms in (3.1) gives only an upper bound. The dimension-dependent error term $\delta^{2}\sqrt{D}$ is due to the variability of the Gaussian noise $\xi$ and unavoidable in the linear case of truncated SVD estimation [3]. Consequently, the early stopped estimator $\widehat{g}_{\tau}$ has the same error rate as $\widehat{g}_{\tau_{\mathfrak{b}}}$ if the remainder term is small enough. In this case of a-posteriori error control, we say that we can adapt to the balanced oracle. In practice, the noise level $\delta$ needs to be estimated, and the threshold $\kappa$ is allowed to deviate from $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]=\delta^{2}D$.
@@
where $\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ is given in (3.2), and the optimal rate of convergence is attained. The early stopping rule $\tau$ is minimax adaptive for the prediction error over Sobolev-type ellipsoids for all regularities $\mu>0$ with $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$ (see Corollary 6.9).
In Section 7, we transfer the results for the prediction error to the reconstruction error, i.e. the error of the CG estimator. Similarly to the prediction error, the analysis of the reconstruction error relies on a nonasymptotic upper bound given in Proposition 7.1, which depends on the error terms and the residual polynomial.
@@
$\sup_{f:\,\lVert f\rVert_{\mu}\leqslant R}\operatorname{{\mathbb{E}}}\big{[}\lVert\widehat{f}_{\tau_{\mathfrak{b}}}-f^{\dagger}\rVert^{2}\big{]}\lesssim_{\mu,p,C_{A}}\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$
with
$\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)\coloneqq R^{(4p+2)/(4\mu p+2p+1)}\delta^{8\mu p/(4\mu p+2p+1)}.$ (3.5)
@@ 4 The interpolated conjugate gradient method
Although we will not use it, the interpolation follows also naturally by defining $r_{t}$ via penalised minimisation. Let
$r_{k+1,\lambda}\coloneqq\operatorname*{argmin}_{p_{k+1}\in\operatorname{Pol}_{k+1,1}}\Big{(}\lVert\mathsf{p_{k+1}}\rVert_{Y}^{2}+\lambda a_{k+1}(p_{k+1})^{2}\Big{)},$ (4.1)
where $a_{k+1}(p)\in{\mathbb{R}}$ denotes the coefficient in front of $x^{k+1}$ for a polynomial $p(x)$ and $\lambda>0$ is a penalty parameter. At the end of Section 9.3, we prove
$r_{k+1,\lambda}=r_{k}+\alpha_{\lambda}(r_{k+1}-r_{k})\quad\text{with}\quad\alpha_{\lambda}=\frac{R_{k}^{2}-R_{k+1}^{2}}{R_{k}^{2}-R_{k+1}^{2}+\lambda a_{k+1}(r_{k+1})^{2}}\in(0,1].$ (4.2)
Since $\alpha_{\lambda}\downarrow 0$ for $\lambda\uparrow\infty$ and $\alpha_{\lambda}=1$ for $\lambda=0$, we recover $r_{t}$ for $t=k+\alpha$, $k=0,\dots,d-1$, $\alpha\in(0,1)$, as $r_{k+1,\lambda}$ for a suitable penalisation parameter $\lambda=\lambda(t)>0$.
We have extended Definition 3.1 to $t\in[0,d]$, keeping the starting values $\widehat{g}_{0}=\widehat{f}_{0}=0$, $r_{0}\equiv 1$ and the relation $A\widehat{f}_{t}=\widehat{g}_{t}$ for $t\in[0,d]$. Note that $r_{t}$ for $t\in(k,k+1]$, $k=0,\dots,d-1$, is an element of $\operatorname{Pol}_{k+1,1}$.
Lemma 4.3 (Properties of the squared residual norm).
  1. (a)
    For $t=k+\alpha$, $k=0,\dots,d-1$, $\alpha\in[0,1]$, the squared residual norm interpolates nonlinearly:
@@
  2. (b)
    The squared residual norm $R_{t}^{2}$ is continuous and decreasing in $t\in[0,d]$.
  3. (c)
    For $t,s\in[0,d]$, we have
@@
and set $r_{0,<}\coloneqq r_{0}\equiv 1$, $r_{0,>}\equiv 0$. If $x_{1,t}$ is clear from the context, we also write $f_{<}(x)\coloneqq f(x){\bf 1}(x<x_{1,t})$, $\lVert\mathsf{f}\rVert_{Y,<}\coloneqq\lVert\mathsf{f}_{<}\rVert_{Y}$ and $\langle{\mathsf{f}},{\mathsf{g}}\rangle_{Y,<}\coloneqq\langle{\mathsf{f_{<}}},{\mathsf{g_{<}}}\rangle_{Y}$ for any functions $f$ and $g$, analogously for ‘$>$’. In particular, we set $\lVert v\rVert_{<}\coloneqq\lVert 1_{<}(AA^{\top})v\rVert$ and $\lVert w\rVert_{<}\coloneqq\lVert 1_{<}(A^{\top}A)w\rVert$ for any vectors $v\in{\mathbb{R}}^{D}$ and $w\in{\mathbb{R}}^{P}$, where $1_{<}(x)\coloneqq{\bf 1}(x<x_{1,t})$, analogously for ‘$>$’ and $\geqslant$’.
For integer iteration indices, the residual polynomials possess several well-known properties, which are collected in Lemma 9.6. These carry over to the interpolated analogues. We state in the following lemma the most important properties, denoting by $\lVert\cdot\rVert_{\mathrm{spec}}$ the spectral norm of a matrix.
@@
    For $k=1,\dots,d-1$, $r_{k}$ has $k$ real simple zeros $0<x_{1,k}<\dots<x_{k,k}<\lVert AA^{\top}\rVert_{\mathrm{spec}}$. For $k=d$, the zeros are the $d$ distinct squared nonzero singular values $0<x_{1,d}=\tilde{\lambda}_{d}^{2}<\dots<x_{d,d}=\tilde{\lambda}_{1}^{2}=\lVert AA^{\top}\rVert_{\mathrm{spec}}$.
@@
    For fixed $x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}]$, $t\mapsto r_{t,<}(x)$ is continuous and nonincreasing on $[0,d]$.
@@
Using the upper bound, we get directly the following result. It will allow us to control the approximation error $A_{t,\lambda}$ from Definition 5.1.